40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 832     Entry time: Wed Aug 13 20:13:35 2008
Author: Yoichi 
Type: Update 
Category: SUS 
Subject: PRM stand-off is glued 
Steve, Janne, Rob, Bob, Koji, Yoichi

We finally managed to balance the PRM and the stand-off is now glued.

The whole procedure was something like this:

(1) Measure the levelness of the optical table. It was done by a bubble level claiming that
the sensitivity is 60 arcsec (roughly 0.3 mrad).
There was no noticeable tilt along the longitudinal direction of the table.

(2) We put a He-Ne laser on one end of the table. Mounted a QPD on a X-Y-Z stage. Put the QPD very
close to the laser and centered it by moving the QPD.
Then we moved the QPD far away from the laser and centered the beam spot in vertical direction
by changing the tilt of the laser mount.
We then moved the QPD close to the laser again and adjusted the height to center it. By repeating
the centering at two locations (near and far) several times, we aligned the laser beam parallel to
the table.

(3) The PRM suspension tower was put on the other end of the optical table, i.e. far from the laser.
The QPD was moved next to the laser to form an optical lever. The height of the QPD is preserved from
the previous step.


(4) A stand-off was picked by a pair of tweezers. By gently lifting the mirror by the bottom earthquake stops,
the tension of the wire was relieved. Then the stand off was slid in below the guide rod.

(5) Using the microscope, it was confirmed that the wire is in the grooves on both sides.

(6) Without damping, it was too much pain to balance the mirror. So we put spare OSEMs in the suspension and
pulled a long cable from the suspension rack to the clean room with a satellite amp.

(7) It turned out that the pinout of the cable is flipped because of the vacuum feed through. So we asked Ben for help.
He made conversion cables. We also found UR OSEM was not responding. Ben opened the satellite box, and we found an op-amp was burnt.
Probably it was because we connected OSEMs wrongly at first and the LED current driver was shorted. We switched the satellite box
from the PRM one to the BS one. Ben will fix the PRM box.
Bob cleaned up some D-Sub converters for the interface with the clean OSEM pigtails.

(8) While waiting for Ben, we also tried to short the OSEM coils for inductive damping. We saw no noticeable change in the Q.

(9) After the OSEMs were connected to the digital control system, Rob tweaked the damping gains a bit to make it work efficiently.

(10) I pushed the stand-off back and forth to make the reflected beam spot centered on the QPD. I used the PZT buzzer to gently move the stand-off.
For fine tune, just touching it is enough. I found it useful to touch it without clamping the mirror, because if it is clamped, we can easily push
it too hard. When the mirror is freely hanging, once the tip of the buzzer touches the stand-off, the mirror escapes immediately. If the mirror
swings wildly by your touch, you pushed it too hard.

(11) After about an hour of struggle, I was able to level the mirror. We used about 1.5m optical lever arm. A rough calibration tells us that the
beam spot is within 0.6mm of the center of the QPD. So the reflected light is deflected by 0.4mrad. That means the mirror
is rotated by 0.2mrad. The OSEMs should have about 30mrad of actuation range. So this should be fine.

(12) We mixed the Vac Seal epoxy and put it under vacuum for 15min to remove bubbles. Actually 15min was not enough for removing bubbles completely. But
stopped there because we did not want the epoxy to be too stiff.
I dipped a thin copper wire into the epoxy and applied it on the top of the stand-off. I found the epoxy is already not fluid enough, so Steve made
another Vac Seal mixture. This time we put it under vacuum for only 3 min.
I also applied the epoxy to the sides of the stand-off.
While working on this, I accidentally touched the side of the PRM. Now there is a drop of epoxy sitting there (upper left of the attached picture).
We decided not to wipe it out because we did not want to screw up the levelness.

(13) We put an incandescent light about 1m away from the suspension to gently warm up the epoxy but not too much. We will leave it overnight to cure the
epoxy.
Attachment 1: img1.jpg  62 kB  | Hide | Hide all
img1.jpg
ELOG V3.1.3-